# Anticipating Object State Changes

### v1

@@ 4 Experimental Setup
Implementation Details: The proposed state anticipation model (dotted rectangle in Fig. 2) is trained on a single NVIDIA TITAN GPU using the Adam optimizer, a batch size of 32, a learning rate of $41e-$ 4, without any temporal augmentations (clip or frame cropping). Short-term associations between neighboring segments of an input video are represented using the pre-extracted SlowFast frame level features from Ego4D. Regarding the selection of the pre- and the post-state keyframes that are introduced in the object recognition model, we exploit the PNR annotations of Ego4D, which correspond to the first frame in each clip when the state change/transition is visible.
We should note that in real-world inference, the action and state lexical histories in the proposed anticipation model will be populated by existing action recognition and object state estimation models trained on respective data of the task.
No OSC
activate
deactivate
construct
deconstruct
deposit
remove
deform
other
Train 2066 4017 1492 4186 1773 14984 15338 4400 15667
Test 1284 1888 617 2289 966 7613 7608 2149 8715
Table 2: Statistics for the Ego4D-OSCA dataset per object state change class. In total the dataset has 61858 training clips and 31846 testing clips.
5 Experimental Results
5.1 Object State Change Anticipation
@@
In Table 4, we compare variants of an object state anticipation model, to highlight the impact of incorporating lexical histories of past actions and object states on object state anticipation performance. Starting with the vision-only model (VID-A), which solely relies on the visual representation of the current action, we observe relatively modest performance levels. However, when lexical histories of past actions are introduced through an oracle recognition model (VNLP (O-Action)), we observe a slight performance improvement, indicating the potential benefit of contextual action information. Notably, incorporating lexical histories of past object states from an oracle recognition model (VNLP (O-State)) leads to significant performance gains, that highlight the importance of considering object state dynamics in anticipation tasks. Further improvements are observed when both types of lexical histories are integrated into the model (VNLP (O-Action, O-State)), demonstrating the synergistic effect of leveraging contextual information from both actions and object states. Overall, the low anticipation scores underscore the inherently challenging nature of the task and the intricacy of the scenarios of the dataset that pose significant challenges for anticipation models.
We also performed experiments in which the oracle-based state estimation was replaced with an actual object state estimator. To fulfill this requirement in our experiments, we adopted the object state classification method of [13], with minor adjustments that we elucidate below. The state history of the non-oracle model relies on the outputs of two distinct state classifiers. Each classifier receives either the first or the last frame of each video segment as input to predict the object state label for the respective frame. The prediction of the state change label for the video segment considers both outputs and is derived based on the following rules. If the object state predictions are $pre\_X$ and $post\_X$, respectively, the inferred state change for the video segment is denoted as $X$. Conversely, if the classifiers predict $pre\_X$ and $post\_Y$, where $X$ and $Y$ are distinct and represent inverse state changes, it is concluded that no state change has occurred. Finally, if neither of the conditions mentioned above is met, the state change prediction defaults to the output of the second classifier; i.e., if the prediction is $post\_Y$, the state change for the video segment is identified as Y. For example, if the predictions of the two classifiers are $pre\_activate$ and $post\_activate$ the prediction for the state change would be $activate$. Likewise, if the predictions are $pre\_activate$ and $post\_deactivate$ the prediction of the state change would be that of no change.
When we replace the oracle object state recognition model with the above-described object state estimator, which showcased a mean frame-wise object state recognition accuracy of $24.1\%$ in this dataset, we observe a significant drop in accuracy (rows 5 & 6 of Table 4). This difference in accuracy underscores the critical role of precise object state recognition in effectively anticipating future object states within dynamic environments.
@@
VNLP (O-Action, State [13]) 28.56 / 92.24%
@@
Table 4: Object state change anticipation model robustness to recognizer performance variability.
5.2 The Impact of Object State and Action Recognition Accuracy
@@ 6 Conclusions
We introduced the new problem of object next-state anticipation and proposed a baseline framework to address it. Our framework is a novel effort in this domain, shedding light on the potential of leveraging contextual information to enhance object state anticipation performance within vision-based models. By integrating lexical histories of past actions and object states with visual information of the recent past, our framework demonstrates notable improvements in anticipation accuracy, underscoring the importance of contextual understanding in dynamic environments. To validate the efficacy of our approach, we adapted a recent, large-scale egocentric dataset, Ego4D [14], introducing a specialized subset for object state anticipation, referred to as Ego4D-OSCA.
An integral aspect of the proposed method involved employing a basic NLP model to handle the lexical histories of previous actions and object states. While this approach offers certain advantages, as discussed in Section 3, it is constrained by its reliance on simplistic lexical structures adhering to predefined structural protocols and language formats. We can substitute these models with LLMs to address these limitations, particularly in capturing nuanced semantics and enhancing adaptability across diverse domains and languages. However, this comes at the extra cost of increased computational and storage demands. Finally, another interesting future direction we plan to explore is studying the problem under zero-shot settings. This could lead to the possibility of developing a method that can anticipate changes that are not included in the training data, or anticipate changes that affect novel classes of objects, or occur as the effect of novel classes of actions.
7 Acknowledgements

### v2

@@ 5 Implementation, Experiments and Results
Implementation Details: The proposed state anticipation model (dotted rectangle in Fig. 6) is trained on a single NVIDIA TITAN GPU using the Adam optimizer, a batch size of 32, a learning rate of $41e-$ 4, without any temporal augmentations (clip or frame cropping). Short-term associations between neighboring segments of an input video are represented using the pre-extracted SlowFast frame level features from Ego4D. Regarding the selection of the pre- and the post-state keyframes that are introduced in the object recognition model, we exploit the PNR annotations of Ego4D, which correspond to the first frame in each clip when the state change/transition is visible. We should note that in real-world inference, the action and state lexical histories in the proposed anticipation model will be populated by existing action recognition and object state estimation models trained on the respective data of the task.
Evaluation metrics: The evaluation of all examined models was conducted using top-1 and top-5 mean accuracy, following standard practices in the relevant literature.
5.1 OSCA Results
In Table 2, we compare variants of an object state anticipation model to highlight the impact of incorporating lexical histories of past actions and object states on the anticipation performance. The vision-only model (VID-A) only relies on the visual representation of the current action. We observe modest performance levels.
@@ 5.1 OSCA Results
OSCA under ideal action and state recognition: When ground-truth lexical histories of past actions are introduced through an oracle recognition model (VNLP (O-Action)), a slight performance improvement indicates the potential benefit of contextual action information. Notably, incorporating lexical histories of past object states from an oracle recognition model (VNLP (O-State)) leads to significant performance gains, that highlight the importance of considering object state dynamics in anticipation tasks. Further improvements are observed when both lexical histories are integrated into the model (VNLP (O-Action, O-State)), demonstrating the synergistic effect of leveraging contextual information from actions and object states. Overall, the low anticipation scores highlight the inherently challenging nature of the task and the intricacy of the dataset scenarios that pose significant challenges for anticipation models.
OSCA under actual action and state recognition: We conducted additional experiments where the oracle-based modules, namely the action recognizer and the current object-agnostic state estimator, were substituted with baseline models. For this experimental setting, the output of these models is utilized to populate the action and state histories, therefore aiding our OSCA model in estimating the anticipated state. We employed the following models: (a) the well-established SlowFast [12] model for action recognition; (b) the object-agnostic state classification method proposed by [17] as the object-agnostic current state classifier, with minor modifications, as described below.
Action recognizer: Regarding the action classification module, we fine-tuned the SlowFast model [12] on a subset of the original Ego4D dataset. Specifically, we obtained the train/validation/test splits using the training set provided for the (LTA) long-term anticipation task of Ego4D based on the $60/20/20$ split scheme. The adaptation of the LTA data into the action recognition task resulted in $5754$ action classes and a total of $\approx 65K$ video clips (with a mean of $\approx 10.7$ sample clips per action class). SlowFast achieved $12.86\%$ Top-1 and $33.69\%$ Top-5 accuracy for the task of current action recognition. This low performance can be attributed to the extensive number of action classes and limited samples per class, as well as in Ego4D’s inherent motion and appearance similarities across different actions, e.g. take cup - take bottle, tie string - tie rope.
@@
VNLP (Action [12] , State [17]) 28.97 / 92.25%
@@
Object state recognizer: Additionally, for the object-agnostic state history, we adapt the model of [17]. This model relies on the outputs of two distinct state classifiers. Each classifier receives the first (pre) or the last (post) frame of each video segment as input to predict the object state label for the respective frame. The prediction of the state change label for the video segment considers both outputs and is derived based on the following rules. If the object state predictions are $pre\_X$ and $post\_X$, respectively, the inferred state change for the video segment is denoted as $X$. Conversely, if the classifiers predict $pre\_X$ and $post\_Y$, where $X$ and $Y$ are distinct and represent inverse state changes, it is concluded that no state change has occurred. Finally, if neither of the conditions mentioned above is met, the state change prediction defaults to the output of the second classifier; i.e., if the prediction is $post\_Y$, the state change for the video segment is identified as Y. For example, if the predictions of the two classifiers are $pre\_activate$ and $post\_activate$ the prediction for the state change would be $activate$. Likewise, if the predictions are $pre\_activate$ and $post\_deactivate$ the prediction of the state change would be that of no change. This object-agnostic state recognizer showcased a mean frame-wise state recognition accuracy of $25.4\%$.
In our experiments when we replaced the oracles with the aforementioned realistic recognition models to populate the action and state history buffers that are being considered by the baseline OSCA model, we observed a significant drop in accuracy (rows 4, 6, & 8 of Table 2). This difference in accuracy underscores the critical role of precise current action and object state recognition in effectively anticipating future object states within dynamic environments.
@@
Table 3: The robustness of the object state change anticipation model is tested to the recognizer performance variability.
5.2 Object State & Action Recognition Impact
@@ 6 Conclusions
This paper introduced the new problem of object state change anticipation during procedural activities. We proposed a novel framework that integrates lexical histories of past actions and object states with recent visual information to enhance anticipation accuracy in vision-based models. By fusing long-term semantic and recent visual information, our framework demonstrates notable improvements in anticipation accuracy, underscoring the importance of contextual understanding in dynamic environments. To validate our approach, we augmented the Ego4D dataset [18] forming a specialized subset noted Ego4D-OSCA. Future work will explore the applicability of LLMs as a replacement for the NLP processing component of the proposed framework, for leveraging their enhanced semantic understanding and in-context learning abilities. We also plan to explore zero-shot settings to enable anticipation of state changes involving novel, previously unseen, objects or actions.
Acknowledgements